OpenStudy (anonymous):

Solve the system using substitution. y=-2x 5x+y=3

OpenStudy (a1234):

y=-2x 5x+y=3 5x - 2x = 3 3x = 3 x = 1 y = -2(1) y = -2 The answer is (1, -2).